OpenStudy (anonymous):

Factor Completely: 4x2 − 81 (4x − 9)(x + 9) (2x + 9)(2x + 9) (2x + 9)(2x − 9) (2x − 9)(2x − 9)

OpenStudy (anonymous):

the 3rd one

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

4 =2^2 x^2= x times x 81= 9^2 Apply it to the rule (a^2-b^2) = (a+b)(a-b)